OpenStudy (anonymous):

Solve the Inequality by using the Test Point Method 3/2x+5 - 1/x-1 ≤ 0

hero (hero):

1. First acknowledge that \(x \ne -\frac{5}{2} \) and \(x \ne 1\) -5/2 and 1 are your first 2 test points. Solve this inequality to get the third test point: \(\frac{3}{2x+5} \le \frac{1}{x-1}\)

hero (hero):

By cross multiplying, you get 3(x-1) ≤ 2x + 5 3x - 3 ≤ 2x + 5 x ≤ 8 Which is the third test point.

hero (hero):

Place all the test points on the number line: 

hero (hero):

We know that x ≤ 8, but x cannot include 1, so we shade that segment. Since The shaded parts cannot overlap, we know the region between -5/2 and 1 does not include the solution. Therefore, the other shaded portion must be x < -5/2